if bob had 4,088 marbles and lost 3/8 of them how much would he have left?
Answer by checkley77(12569)

Thank you very much but why times it by 5/ 8 could you please show me how you
got this trying to learn.... again thanks so very much
The reason for the multiplication of 5/8 is because (1-3/8=5/8).
If 3/8 are lost then only 5/8 remain & you need to know how many remain so I changed the 3/8 being lost to the 5/8 that remained & did the necessary multiplication below.
4,088*5/8=2,555 marbles are left.